OpenStudy (anonymous):

x^{3/8} - 12 x^{3/16} + 36 = 0

OpenStudy (anonymous):

Hello there. Any ideas?

hartnn (hartnn):

can u solve quadratic equations ?

OpenStudy (anonymous):

Yeah, but we're supposed to use substitution to solve this one.

hartnn (hartnn):

right so substitute x^{3/16}=y

OpenStudy (anonymous):

Right, I've tried that. And ended up with y^2 - 12y + 36, which I've then reduced to (y-6)^2=0

OpenStudy (anonymous):

But apparently 6 is wrong....

hartnn (hartnn):

because y=6 x is NOT 6

hartnn (hartnn):

u need to find x from x^{3/16}=y

OpenStudy (anonymous):

Right, right. So I just find what 6^(3/16) equals?

hartnn (hartnn):

nope

OpenStudy (anonymous):

Then what should I do?

hartnn (hartnn):

x^{3/16}=y x= y^{16/3} x= 6^{16/3}

hartnn (hartnn):

got that ^^ ?

OpenStudy (anonymous):

Ohhh....okay, yeah. Thank you so much!

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Thanks for the medal as well! ^_^

OpenStudy (anonymous):

Have a good night.

hartnn (hartnn):

good night to u, and good morning to me :)

OpenStudy (anonymous):

Yes, I suppose so xp

OpenStudy (anonymous):

Actually, wait.